# Equivariant and Invariant Parametrized Topological Complexity

### v3

@@ 4.3. Invariance Theorem
[…]
@@
Suppose $U$ is a $(G\times G)$-invariant open subset of $E\times_{B/G}E$ with a $(G\times G)$-equivariant section $s$ of $\Psi$ over $U$. Then $\overline{U}:=(\pi_{E}\times\pi_{E})(U)$ is an open subset of $\overline{E}\times_{\overline{B}}\overline{E}$, by Lemma 4.27. As $U$ is $(G\times G)$-invariant, it follows $U$ is saturated with respect to $\pi_{E}\times\pi_{E}$. Hence, $\pi_{E}\times\pi_{E}\colon U\to\overline{U}$ is a quotient map. Then, by universal property of quotient maps, there exists a unique continuous map $\overline{s}\colon\overline{U}\to E^{I}_{B}$ such that the following diagram
[…]
@@
Suppose $\overline{U}$ is an open subset of $\overline{E}\times_{\overline{B}}\overline{E}$ with section $\overline{s}$ of $\overline{\Pi}$ over $\overline{U}$. Then, by Theorem 3.2 for the trivial group action, there exists a homotopy $\overline{H}\colon\overline{U}\times I\to\overline{E}\times_{\overline{B}}\overline{E}$ such that $\overline{H}_{0}$ is the inclusion map of $i_{\overline{U}}\colon\overline{U}\hookrightarrow\overline{E}\times_{\overline{B}}\overline{E}$ and $\overline{H}_{1}$ takes value in $\Delta(\overline{E})$.
@@
[…]
commutes. As the $G$-action on $E$ is free, it follows the action of $G\times G$ on $E\times_{E/G}E$ (and $U$) is free. Hence, by the Covering Homotopy Theorem of Palais [7, Theorem II.7.3] and Lemma 4.27, it follows there exists a $(G\times G)$-homotopy $H\colon U\times I\to E\times_{B/G}E$ such that $H_{0}=i_{U}\colon U\hookrightarrow E\times_{B/G}E$ and $(\pi_{E}\times\pi_{E})\circ H=\overline{H}\circ((\pi_{E}\times\pi_{E})\times\mathrm{id}_{I})$. As $\overline{H}_{1}$ takes value in $\Delta(\overline{E})$, it follows $H_{1}$ takes values in $E\times_{E/G}E$. Hence, by Theorem 4.7, we get a $(G\times G)$-equivariant section of $\Psi$ over $U$. Thus, $\mathrm{TC}^{G}[p\colon E\rightarrow B]\leq\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]$. ∎
We note that the main theorem in Lubawski and Marzantowicz’s paper, as stated in Theorem 2.25, can be recovered from Theorem 4.29 by taking the base space $B$ to be a point. Now, we state some corollaries of this theorem.
@@
Observe that $G$ acts freely on $E$. Hence, the result follows from Corollary 4.14 and Theorem 4.29. ∎
@@
By Proposition 4.25, we have $\overline{p}\colon\overline{E}\rightarrow\overline{B}$ is a locally trivial fibration with fibre $F$. We note that the fibre of the parametrized fibration $\overline{\Pi}\colon(\overline{E})^{I}_{\overline{B}}\to\overline{E}\times_{\overline{B}}\overline{E}$ induced by $\overline{p}\colon\overline{E}\to\overline{B}$ is the loop space $\Omega F$, which is contractible since $F$ is contractible. Hence, $\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]=1$ by obstruction theory. Thus, the result follows from the Invariance Theorem 4.29. ∎
@@ 5. Computations for equivariant Fadell-Neuwirth fibrations
$F(\mathbb{R}^{d},s):=\{(x_{1},\dots,x_{s})\in(\mathbb{R}^{d})^{s}\mid x_{i}\neq x_{j}~{}\text{ for }~{}i\neq j\}.$
Definition 5.1 ([15]).
@@
The space $F(\mathbb{R}^{d},s+t)$ admits an action of the permutation group $\Sigma_{s}$ defined as follows. For $\sigma\in\Sigma_{s}$, let
@@
Similarly, $\Sigma_{s}$ acts on $F(\mathbb{R}^{d},s)$ by permuting the coordinates $(x_{1},\dots,x_{s})$. Notice that the map $p$ in Definition 5.1 is $\Sigma_{s}$-equivariant. In fact, in [12], it was demonstrated that this map is a $\Sigma_{s}$-fibration.
For the rest of the section, we will use the notation $p\colon E\to B$ for the equivariant Fadell-Neuwirth fibration. The fibre $F$ of $p$ is the configuration space of $t$ points on $\mathbb{R}^{d}\setminus\mathcal{O}_{s}$, where $\mathcal{O}_{s}$ represents the configuration of $s$ distinct points in $\mathbb{R}^{d}$. More precisely, $F=F(\mathbb{R}^{d}\setminus\mathcal{O}_{s},t)$.
@@
[…]
@@
$\mathrm{TC}^{\Sigma_{1}}[p\colon F(\mathbb{R}^{d},1+t)\to F(\mathbb{R}^{d},1)]$ $=\mathrm{TC}[p\colon F(\mathbb{R}^{d},1+t)\to F(\mathbb{R}^{d},1)]$ by Proposition 4.4 (1)
$=\mathrm{TC}(F(\mathbb{R}^{d}\setminus\mathcal{O}_{1},t))$ by [8, Example 4.2]
@@
since $F(\mathbb{R}^{d}\setminus\mathcal{O}_{1},t)$ is homotopy equivalent to $F(\mathbb{R}^{d},t+1)$. We note that topological complexity of configuration spaces was computed by Farber and Grant in [17]. Thus, we will assume that $s\geq 2$.
The parametrized topological complexity of these fibrations were computed in [8] and [9]. In particular, they proved the following result:
Theorem 5.3 ([8, Theorem 9.1] and [9, Theorem 4.1]).
Suppose $s\geq 2$, $t\geq 1$. Then
@@
We will now demonstrate that the invariant parametrized topological complexity of the Fadell–Neuwirth fibrations coincides with that of the corresponding orbit fibrations. Furthermore, it is bounded below by the parametrized topological complexity of the Fadell–Neuwirth fibrations. This implies that the complexity of the universal motion planning algorithm is greater when the order in which obstacles (mines) are placed is irrelevant. This is something one would expect, since the motion planners, in a sense, satisfies an extra condition, i.e., they remain unchanged under the reordering of the obstacles (mines).
@@
We note that the equivariant Fadell-Neuwirth fibrations are locally $\Sigma_{s}$-trivial with $\Sigma_{s}$ acting trivially on the fibre $F$, see [12, Section 5.1]. Since $F(\mathbb{R}^{d},s)$ is a manifold, it is paracompact Hausdorff. Hence, by Proposition 4.25, it follows that the induced map $\overline{p}$ is a locally trivial fibration with fibre $F$.
As the action of $\Sigma_{s}$ on $F(\mathbb{R}^{d},s+t)$ and $F(\mathbb{R}^{d},s)$ is free, and since $F(\mathbb{R}^{d},s+t)$ and $F(\mathbb{R}^{d},s)$ are manifolds, it follows that $\overline{F(\mathbb{R}^{d},s+t)}$ and $\overline{F(\mathbb{R}^{d},s)}$ are also manifolds. Thus, the result follows from Corollary 4.30. ∎
@@
Suppose $p\colon E\to B$ denotes the equivariant Fadell-Neuwirth fibration with fibre $F$. Then
  1. (1)
    the space $E\times_{B/G}E$ is $(d-2)$-connected, and
  2. (2)
    $\dim\left(\overline{E}\times_{\overline{B}}\overline{E}\right)=\dim(B)+2\dim(F)=ds+2dt$.
  3. (3)
    $\mathrm{hdim}(\overline{E}\times_{\overline{B}}\overline{E})\leq\mathrm{hdim}(\overline{B})+2\dim(F)=(d-1)(s-1)+2dt$.
@@
(1) Since $\Sigma_{s}$ is a finite group acting freely on a Hausdorff space $B$, it follows $\pi_{B}\colon B\to B/G$ is a covering map. Hence, $\pi_{B}$ is a fibration. Thus, $\pi_{1}\colon E\times_{B/G}E\to E$ is a fibration with fibre $\coprod_{g\in G}F$ since the following diagram
${E\times_{B/G}E}$${E}$${E}$${B/G}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{\pi_{B}\circ p}$$\scriptstyle{\pi_{B}\circ p}$
is a pullback. As $E$ and $F$ are $(d-2)$-connected (see discussion after the statement of Theorem 4.1 in [9]), it follows that the space $E\times_{B/G}E$ is $(d-2)$-connected.
(2) As $\overline{p}\colon\overline{E}\to\overline{B}$ is a locally trivial fibration with fibre $F$, it follows that the obvious map $\overline{E}\times_{\overline{B}}\overline{E}\to\overline{B}$ is a locally trivial fibration with fibre $F\times F$. Hence,
$\dim(\overline{E}\times_{\overline{B}}\overline{E})=\dim(F\times F)+\dim(\overline{B})=2\dim(F)+\dim(\overline{B}).$
Since $\Sigma_{s}$-action on the manifold $B$ is free, we get that $\overline{B}$ is a manifold with $\dim(\overline{B})=\dim(B)$.
(3) The homotopy dimension of the unordered configuration space $\overline{B}$ is $(d-1)(s-1)$ (see [2, Example 7.1.12] and [6, Theorem 3.13]). Hence, by Proposition 4.26, the claim follows. ∎
We would like to thank Professor Jesus Gonzalez for pointing us to an appropriate reference on the equivariant $\mathrm{CW}$-complex structure on the ordered configuration space, from which we can determine the homotopy dimension of the unordered configuration space. This result is used in the following theorem and Theorem 5.11.
@@
$\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R}^{d},s+t)\to F(\mathbb{R}^{d},s)]=\begin{cases}2t+s,&\text{if $d$ is odd},\\ \text{either }2t+s-1\text{ or }2t+s&\text{if $d$ is even}.\end{cases}$ (7)
@@
established in Theorem 5.4, together with Theorem 5.3 yields the desired lower bound.
@@
if and only if $(2t+s)$-fold join
@@
of the fibration $\overline{\Pi}$ admits a global section, see [32, Theorem 3]. Note that the loop space $\Omega F$ is path-connected, since $\Omega F$ is $(d-3)$-connected and $d\geq 3$. Therefore, $*_{2t+s}\Omega F$ is simply connected, and hence $k$-simple for all $k$. Thus, the obstruction to a global section of $\overline{\Pi}_{2t+s}$ lies in the cohomology classes
@@
see [31, Theorem 34.2]. As $*_{2t+s}\Omega F$ is $((2t+s)(d-1)-2)$-connected, we have $H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F))=0$ for $k\leq(2t+s)(d-1)-2$.
@@
Consider the Serre spectral sequence $E^{p,q}_{r}$ with local coefficients for the fibration (8), see [30, Theorem 2.9]. Then $E_{2}^{p,q}=H^{p}(\overline{B},H^{q}(F\times F,\mathcal{M}))$, and the spectral sequence converges to $H^{p+q}(\overline{E}\times_{\overline{B}}\overline{E},\mathcal{M})$. If either $p>\mathrm{hdim}(\overline{B})$ or $q>\mathrm{hdim}(F\times F)$, then $H^{p}(\overline{B},H^{q}(F\times F,\mathcal{M}))=0$. Hence,

### v4

@@ 4.3. Invariance Theorem
$E^{I}_{B}\times_{E/G}E^{I}_{B}$$E\times_{B/G}E$$(\overline{E})^{I}_{\overline{B}}$$\overline{E}\times_{\overline{B}}\overline{E},$$\scriptstyle{\Psi}$$\scriptstyle{f}$$\scriptstyle{\pi_{E}\times\pi_{E}}$$\scriptstyle{\overline{\Pi}}$
@@
Suppose $U$ is a $(G\times G)$-invariant open subset of $E\times_{B/G}E$ with a $(G\times G)$-equivariant section $s$ of $\Psi$ over $U$. Then $\overline{U}:=(\pi_{E}\times\pi_{E})(U)$ is an open subset of $\overline{E}\times_{\overline{B}}\overline{E}$, by Lemma˜4.24. As $U$ is $(G\times G)$-invariant, it follows that $U$ is saturated with respect to $\pi_{E}\times\pi_{E}$. Hence, $\pi_{E}\times\pi_{E}\colon U\to\overline{U}$ is a quotient map. Then, by the universal property of quotient maps, there exists a unique continuous map $\overline{s}\colon\overline{U}\to E^{I}_{B}$ such that the following diagram
${\overline{E}^{I}_{B}}U¯{\overline{U}}$$\scriptstyle{f\circ s}$$\scriptstyle{\pi_{E}\times\pi_{E}}$$\scriptstyle{\overline{s}}$
@@
Recall that a space $X$ is called hereditary paracompact if every subspace of $X$ is paracompact. Equivalently, every open subspace of $X$ is paracompact.
@@
We note that, in Theorem˜2.10, it was established that $\overline{p}$ is a fibration. Suppose $\overline{U}$ is an open subset of $\overline{E}\times_{\overline{B}}\overline{E}$ with section $\overline{s}$ of $\overline{\Pi}$ over $\overline{U}$. Then, by Theorem˜3.2 for the trivial group action, there exists a homotopy $\overline{H}\colon\overline{U}\times I\to\overline{E}\times_{\overline{B}}\overline{E}$ such that $\overline{H}_{0}$ is the inclusion map of $i_{\overline{U}}\colon\overline{U}\hookrightarrow\overline{E}\times_{\overline{B}}\overline{E}$ and $\overline{H}_{1}$ takes values in $\Delta(\overline{E})$.
@@
$U\times\{0\}$$E\times_{B/G}E$$U\times I$$\overline{U}\times I$$\overline{E}\times_{\overline{B}}\overline{E}$$\scriptstyle{\pi_{E}\times\pi_{E}}$$\scriptstyle{(\pi_{E}\times\pi_{E})\times\mathrm{id}_{I}}$$\scriptstyle{\overline{H}}$
commutes. As the $G$-action on $E$ is free, it follows the action of $G\times G$ on $E\times_{B/G}E$ and $U$ is free. Hence, the homotopy $\overline{H}$ preserves the orbit structure. Since $G$ is a compact Lie group, by the Covering Homotopy Theorem of Palais [bredon-transformation-groups, Theorem II.7.3] and Lemma˜4.24, it follows that there exists a $(G\times G)$-homotopy $H\colon U\times I\to E\times_{B/G}E$ such that $H_{0}=i_{U}\colon U\hookrightarrow E\times_{B/G}E$ and $(\pi_{E}\times\pi_{E})\circ H=\overline{H}\circ((\pi_{E}\times\pi_{E})\times\mathrm{id}_{I})$. As $\overline{H}_{1}$ takes value in $\Delta(\overline{E})$, it follows $H_{1}$ takes values in $E\times_{E/G}E$. Hence, by Theorem˜4.6, we get a $(G\times G)$-equivariant section of $\Psi$ over $U$. Thus, $\mathrm{TC}^{G}[p\colon E\rightarrow B]\leq\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]$. ∎
We note that the main theorem in Lubawski and Marzantowicz’s paper, as stated in Theorem˜2.31, can be recovered from Theorem˜4.26 by taking the base space $B$ to be a point. Now, we state some corollaries of this theorem.
@@
Observe that $G$ acts freely on $E$ as well. Hence, the result follows from Corollary˜4.13 and Theorem˜4.26. ∎
@@
By Proposition˜4.22, we have $\overline{p}\colon\overline{E}\rightarrow\overline{B}$ is a locally trivial fibration with fibre $F$. We note that the fibre of the parametrized fibration $\overline{\Pi}\colon(\overline{E})^{I}_{\overline{B}}\to\overline{E}\times_{\overline{B}}\overline{E}$ induced by $\overline{p}\colon\overline{E}\to\overline{B}$ is the loop space $\Omega F$, which is contractible since $F$ is contractible. Hence, $\mathrm{TC}[\overline{p}\colon\overline{E}\rightarrow\overline{B}]=1$ by obstruction theory. Thus, the result follows from the Invariance Theorem˜4.26. ∎
Remark 4.29.
Suppose $p\colon B\times F\to B$ is the trivial $G$-fibration. If $G$ acts trivially on $F$, then in Proposition˜4.5, we showed that
$\mathrm{TC}^{G}[p\colon B\times F\to B]=\mathrm{TC}(F).$
In general, however, if $G$ acts non-trivially on $F$, then the following inequality need not hold:
$\mathrm{TC}^{G}[p\colon B\times F\to B]=\mathrm{TC}^{G}(F).$
For example, let $E=S^{1}\times S^{1}$ and $B=S^{1}$. If $G=S^{1}$ acts on $B$ by left multiplication and diagonally on $E$, then
$\mathrm{TC}^{S^{1}}[p\colon S^{1}\times S^{1}\to S^{1}]$ $=\mathrm{TC}[p/S^{1}\colon(S^{1}\times S^{1})/S^{1}\to S^{1}/S^{1}]$ by Theorem˜4.26
$=\mathrm{TC}[S^{1}\to\{*\}]$
$=\mathrm{TC}(S^{1})$
$=2.$
But $\mathrm{TC}^{S^{1}}(S^{1})=\mathrm{TC}(\{*\})=1$ by Theorem˜2.31.
Example 4.30.
Suppose $G$ is a compact Lie group. Let $\tau\colon P\to B$ be a principal $G$-bundle such that $\overline{P}\times\overline{P}$ is hereditary paracompact and $B$ is paracompact. Let $X$ be a path-connected $G$-space, and $q\colon P\times_{G}X\to B$ be the associated fibre bundle with fibre $X$ and structure group $G$. Then, by [Sequential-PTC-and-related-invariants, Theorem 3.4], it follows that
$\mathrm{TC}(X)\leq\mathrm{TC}[q\colon P\times_{G}X\to B]\leq\mathrm{TC}_{G}(X),$
where $p\colon P\times X\to P$ is the trivial $G$-fibration. Moreover, we have
$\mathrm{TC}^{G}[p\colon P\times X\to P]=\mathrm{TC}[q\colon P\times_{G}X\to B]$
by Theorem˜4.26.
Example 4.31.
Now we list some examples of $G$-spaces $X$ for which $\mathrm{TC}(X)=\mathrm{TC}_{G}(X)$. Consequently, for each these examples and for any principal $G$-bundles $\tau\colon P\to B$, one has
$\mathrm{TC}^{G}[p\colon P\times X\to P]=\mathrm{TC}(X),$
where $p\colon P\times X\to P$ is the trivial $G$-fibration.
  1. (1)
    Suppose $n_{1},\dots,n_{m}$ are positive integers. Let $G=\mathbb{Z}_{2}$ act on each $S^{n_{i}}$ via the antipodal action. If $G$ acts diagonally on the product $\prod_{i=1}^{m}S^{n_{i}}$, then by [eqtcprodineq, Corollary 4.4] we have
    $\mathrm{TC}_{G}\left(\prod_{i=1}^{m}S^{n_{i}}\right)=\mathrm{TC}\left(\prod_{i=1}^{m}S^{n_{i}}\right)=m+l+1,$
    where $l$ denotes the number of even-dimensional spheres among the factors.
  2. (2)
    Suppose $n_{1},\dots,n_{m}$ are integers with $n_{i}\geq 2$ for each $i$. Let $G=\mathbb{Z}_{2}$ act on $S^{n_{i}}$ by reflection, defined as multiplication by 1-1 in one of the coordinates. If $G$ acts diagonally on the $\prod_{i=1}^{m}S^{n_{i}}$, then
    $\mathrm{TC}_{G}\left(\prod_{i=1}^{m}S^{n_{i}}\right)$ $\leq\sum_{i=1}^{m}\mathrm{TC}_{G}(S^{n_{i}})-(m-1)$ [eqtcprodineq, Theorem 4.2]
    $=3m-(m-1)$ [colmangranteqtc, Example 5.9]
    $=2m+1.$
    Moreover, it follows from [Rudyak2014, Corollary 3.12] that
    $\mathrm{TC}\left(\prod_{i=1}^{m}S^{n_{i}}\right)=m+l+1,$
    where $l$ denotes the number of even-dimensional spheres among the factors. Hence, if $X=\prod_{i=1}^{m}S^{n_{i}}$ with all $n_{i}$ even, then $\mathrm{TC}_{G}(X)=\mathrm{TC}(X)=2m+1$.
  3. (3)
    Let $F$ be a $2n$-dimensional quasitoric manifold. The $\mathbb{Z}_{2}$-action on $F$ is described in [B-D-S, Equation 3.2], and it is shown in [B-D-S, Corollary 2.10 and Proposition 3.11] that $\mathrm{TC}_{\mathbb{Z}_{2}}(F)=\mathrm{TC}(F)=2n+1$. We refer the reader to [B-D-S] for further details.
  4. (4)
    Let $F=\mathrm{Gr}_{d}(\mathbb{C}^{n})$ be the complex Grassmannian manifold of complex dimension d(nd)d(n-d). Then $F$ admits a $\mathbb{Z}_{2}$-action induced by complex conjugation. It follows from [D-S, Example 6.6] that $\mathrm{TC}_{\mathbb{Z}_{2}}(F)=2d(n-d)+1=\mathrm{TC}(F)$. In particular, for complex projective space $\mathbb{C}P^{n}$, we have $\mathrm{TC}_{\mathbb{Z}_{2}}(\mathbb{C}P^{n})=2n+1=\mathrm{TC}(F)$. We refer the reader to [D-S, Section 5] for further details.
Example 4.32.
Here we list some examples in which the invariant topological complexity of $G$-fibration is trivial.
  1. (1)
    Suppose $G$ is a compact Lie group. Let $\tau\colon P\to B$ be a principal $G$-bundle such that $\overline{P}\times\overline{P}$ is hereditary paracompact and $B$ is paracompact. Then
    $\mathrm{TC}^{G}[p\colon P\to B]=\mathrm{TC}[\overline{p}\colon\overline{P}\to B]=1,$
    since $\overline{p}$ is a homeomorphism.
  2. (2)
    Let $E$ and $B$ be path-connected and locally path-connected spaces such that $\overline{E}\times\overline{E}$ is hereditary paracompact. Suppose $p\colon E\to B$ is a regular covering map and $G$ is its group of covering transformations. If $G$ is a compact Lie group, then
    $\mathrm{TC}^{G}[p\colon E\to B]=\mathrm{TC}[\overline{p}\colon\overline{E}\to B]=1,$
    since $\overline{p}$ is a homeomorphism by [Munkres-Topology, Theorem 81.6].
@@ 5. Computations for equivariant Fadell-Neuwirth fibrations
[…]
Definition 5.1 ([FadellNeuwirth]).
@@
The space $F(\mathbb{R}^{d},s+t)$ admits a free action of the permutation group $\Sigma_{s}$, defined as follows. For $\sigma\in\Sigma_{s}$, let
@@
Similarly, $\Sigma_{s}$ acts freely on $F(\mathbb{R}^{d},s)$ by permuting the coordinates $(x_{1},\dots,x_{s})$. Notice that the map $p$ in Definition˜5.1 is $\Sigma_{s}$-equivariant. In fact, in [D-EqPTC], it was demonstrated that this map is a $\Sigma_{s}$-fibration.
For the rest of the section, we will use the notation $p\colon E\to B$ for the equivariant Fadell-Neuwirth fibration. The fibre $F$ of $p$ is the configuration space of $t$ points on $\mathbb{R}^{d}\setminus\mathcal{O}_{s}$, where $\mathcal{O}_{s}$ denotes a set of $s$ distinct points in $\mathbb{R}^{d}$. More precisely, $F=F(\mathbb{R}^{d}\setminus\mathcal{O}_{s},t)$.
@@
If $t=0$, then $p$ is the identity map. In particular, $p$ is a trivial $\Sigma_{s}$-fibration with fibre a point $\{*\}$. Hence, by Proposition˜4.5, it follows that
@@
$\mathrm{TC}^{\Sigma_{1}}[p\colon F(\mathbb{R}^{d},1+t)\to F(\mathbb{R}^{d},1)]$ $=\mathrm{TC}[p\colon F(\mathbb{R}^{d},1+t)\to F(\mathbb{R}^{d},1)]$ by Proposition˜4.4 (1)
$=\mathrm{TC}(F(\mathbb{R}^{d}\setminus\mathcal{O}_{1},t))$ by [farber-para-tc, Example 4.2]
@@
since $F(\mathbb{R}^{d}\setminus\mathcal{O}_{1},t)$ is homotopy equivalent to $F(\mathbb{R}^{d},t+1)$, see [fadell-husseini-conf-spaces, Page 15]. We note that topological complexity of configuration spaces was computed by Farber and Grant in [TC-of-configuration-spaces]. Thus, we will assume that $s\geq 2$.
If $d=1$, then the fibre $F(\mathbb{R}\setminus\mathcal{O}_{s},t)$ of $p$ is not path-connected. Hence, by Corollary˜4.13, it follows that
$\infty=\mathrm{TC}(F(\mathbb{R}\setminus\mathcal{O}_{s},t))\leq\mathrm{TC}^{\Sigma_{s}}[p\colon F(\mathbb{R},s+t)\to F(\mathbb{R},s)].$
Thus, we will assume that $d\geq 2$.
The parametrized topological complexity of these fibrations was computed in [farber-para-tc] and [PTCcolfree]. In particular, they proved the following result:
Theorem 5.3 ([farber-para-tc, Theorem 9.1] and [PTCcolfree, Theorem 4.1]).
Suppose $s\geq 2$, $t\geq 1$ and $d\geq 2$. Then
@@
We will now demonstrate that the invariant parametrized topological complexity of the Fadell–Neuwirth fibrations coincides with that of the corresponding orbit fibrations. Furthermore, it is bounded below by the parametrized topological complexity of the Fadell–Neuwirth fibrations. Interpreting the parametrized topological complexity of the Fadell–Neuwirth fibration as describing the motion planning problem of $t$ submarines navigating in the presence of $s$ mines whose positions are unknown, and that of the orbit Fadell–Neuwirth fibration as describing the corresponding problem when the order of the mines is irrelevant, the resulting increase in complexity is entirely consistent with intuition. When the order of the mines is irrelevant, the motion planners must satisfy an additional constraint: they are required to remain invariant under any reordering of the mines. In other words, the planner must yield the same motion plan regardless of how the mines are permuted. This symmetry condition imposes additional structural restrictions on the class of admissible planners, thereby increasing the overall complexity of constructing such a universal algorithm.
@@
We note that the equivariant Fadell-Neuwirth fibrations are locally $\Sigma_{s}$-trivial with $\Sigma_{s}$ acting trivially on the fibre $F$, see [D-EqPTC, Section 5.1]. Since $F(\mathbb{R}^{d},s)$ is a manifold, it is paracompact Hausdorff. Hence, by Proposition˜4.22, it follows that the induced map $\overline{p}$ is a locally trivial fibration with fibre $F$.
As the action of $\Sigma_{s}$ on $F(\mathbb{R}^{d},s+t)$ and $F(\mathbb{R}^{d},s)$ is free, and since $F(\mathbb{R}^{d},s+t)$ and $F(\mathbb{R}^{d},s)$ are manifolds, it follows that $\overline{F(\mathbb{R}^{d},s+t)}$ and $\overline{F(\mathbb{R}^{d},s)}$ are also manifolds. Thus, the result follows from Corollary˜4.27. ∎
@@
Suppose $p\colon E\to B$ denotes the equivariant Fadell-Neuwirth fibration with fibre $F$. Then $\mathrm{hdim}(\overline{E}\times_{\overline{B}}\overline{E})\leq(d-1)(s-1)+2dt$.
@@
Since $\overline{p}\colon\overline{E}\to\overline{B}$ is a locally trivial fibration with fibre $F$, it follows that the natural map $\overline{E}\times_{\overline{B}}\overline{E}\to\overline{B}$ is also a locally trivial fibration with fibre $F\times F$. The homotopy dimension of the unordered configuration space $\overline{B}$ is (d1)(s1)(d-1)(s-1) (see [B-Z, Theorem 3.13]). Hence, by Proposition˜4.23, the desired claim follows. ∎
@@
[…]
@@
established in Theorem˜5.4, together with Theorem˜5.3 yields the desired lower bound.
@@
if and only if $(2t+s)$-fold fibrewise join
@@
of the fibration $\overline{\Pi}$ admits a global section, see [Sva, Theorem 3]. Note that the loop space $\Omega F$, which is the fibre of $\Pi$, is path-connected because $\Omega F$ is (d3)(d-3)-connected and $d\geq 3$. Therefore, the fibre $_{2t+s}\Omega F$ of $\overline{\Pi}_{2t+s}$ is simply connected, and hence $k$-simple for all $k$. Thus, the obstructions to a global section of $\overline{\Pi}_{2t+s}$ lie in the cohomology groups
@@
see [Steenrod, Theorem 34.2]. Since $_{2t+s}\Omega F$ is ((2t+s)(d1)2)((2t+s)(d-1)-2)-connected (see [Sva, Theorem 5] or [grant2019symmetrized, Theorem 3.5] for the homotopy connectivity of joins), it follows that $H^{k+1}(\overline{E}\times_{\overline{B}}\overline{E};\pi_{k}(*_{2t+s}\Omega F))=0$ for $k\leq(2t+s)(d-1)-2$.
@@
Consider the Serre spectral sequence $E^{p,q}_{r}$ with local coefficients for the fibration (11), see [Siegel-Spectral-sequence, Theorem 2.9]. Then $E_{2}^{p,q}=H^{p}(\overline{B},H^{q}(F\times F,\mathcal{M}))$, and the spectral sequence converges to $H^{p+q}(\overline{E}\times_{\overline{B}}\overline{E},\mathcal{M})$. If $p>\mathrm{hdim}(\overline{B})$ or $q>\mathrm{hdim}(F\times F)$, then $H^{p}(\overline{B},H^{q}(F\times F,\mathcal{M}))=0$. Hence,